# Algebraic and geometric properties of homeomorphism groups of ordinals

### v1

@@ 1 Introduction
Second, if $E=\omega^{\alpha+1}+1$ for some countable ordinal $\alpha$, then our first slate of results fail for $\operatorname{MCG}(\mathbb{S}^{2}\smallsetminus E)$. Let us explain: Using the work of Domat–Dickmann [11], Malestein–Tao [22] showed that $\operatorname{MCG}(\mathbb{S}^{2}\smallsetminus(\omega+1))$ surjects onto the additive group of real numbers. Now, there is an epimorphism $\operatorname{MCG}(\mathbb{S}^{2}\smallsetminus E)\to\operatorname{MCG}(\mathbb{S}^{2}\smallsetminus(\omega+1))$ obtained by “filling in” the ends that are neither maximal nor next-to-maximal rank (this is called a forgetful homomorphism in the literature). It follows that $\operatorname{MCG}(\mathbb{S}^{2}\smallsetminus E)$ also surjects onto the reals; in particular, $\operatorname{MCG}(\mathbb{S}^{2}\smallsetminus E)$ is neither uniformly perfect, finitely normally generated, strongly bounded, nor strongly distorted. Therefore, our results together with the work of Malestein–Tao, show that the failure of $\operatorname{Homeo}(\mathbb{S}^{2}\smallsetminus E)$ to be strongly bounded, uniformly perfect, and finitely normally generated, stems from a two-dimensional obstruction, i.e., it cannot be detected by the action on the end space. It is also interesting to note that the results only fail in the category of abstract groups, meaning their analogs in the category of topological groups hold; in particular, Mann–Rafi [23] showed that $\operatorname{MCG}(\mathbb{S}^{2}\smallsetminus E)$ is coarsely bounded, and it follows from the work of the last author with Lanier [20] that $\operatorname{MCG}(\mathbb{S}^{2}\smallsetminus E)$ is topologically normally generated by a single mapping class (Baik [4] also showed finite topological normal generation for $\operatorname{MCG}(\mathbb{S}^{2}\smallsetminus(\omega^{\alpha+1}\cdot d+1))$, where $\alpha$ is a countable ordinal and $d\in\mathbb{N}$).
@@ Acknowledgments
The authors thank Kathryn Mann for the reference to Galvin’s work and Ferrán Valdez for a discussion of his forthcoming work with collaborators on the geometry of homeomorphism groups of ordinals.
@@ 2 Ordinals and their topology
We provide an introduction to ordinals and their topology, aimed at non-experts. In the first three subsections we cover the basics of ordinals, following [10, §4.1–2] and [18, Chapter 2].
@@ 2.1 Defining the ordinals
We will use von Neumman’s definition of the ordinals.
@@
A set $\alpha$ is an ordinal if it satisfies the following:
@@
Every element of an ordinal is itself an ordinal. An ordinal can be well-ordered by set inclusion, i.e., if $\alpha$ is an ordinal and $\beta,\gamma\in\alpha$, declaring $\beta\leq\gamma$ if and only if $\beta\subseteq\gamma$ yields a well-ordering relation on $\alpha$. Note that if $\beta,\gamma\in\alpha$, then $\gamma<\beta$ if and only if $\gamma\in\beta$. In particular, if $\beta\in\alpha$, then $\beta=\{\eta\in\alpha:\eta<\beta\}$. If $\alpha$ and $\beta$ are order-isomorphic ordinals, then $\alpha=\beta$; it now follows from basic properties of well-ordered sets that given two ordinals $\alpha$ and $\beta$, either $\alpha=\beta$, $\alpha\in\beta$, or $\beta\in\alpha$. We can therefore compare any two distinct ordinals, writing $\beta<\alpha$ if and only if $\beta\in\alpha$ (here, we have slightly abused notation, using $<$ to denote both a relation on a given ordinal and as a comparison of two distinct ordinals; however, no confusion arises with this abuse). In other words,
The ordinal $\alpha$ is the set of all ordinals less than $\alpha$.
[…]
@@
Let $\alpha$ be an ordinal, and let $S\subset\alpha$. Suppose $S$ satisfies the following: if $\lambda\in\alpha$ and $\beta\in S$ for all $\beta<\lambda$, then $\lambda\in S$. Then $S=\alpha$. ∎
The above statement is a bit awkward because we are avoiding using the languages of classes, allowing us to make a statement that is valid in ZFC.
@@ 2.2 Ordinal arithmetic
With the definition of an ordinal, it follows from the basic theory of well-ordered sets that every well-ordered set is order isomorphic to a unique ordinal, a fact we will use to define ordinal arithmetic.
@@
Let $\alpha$ and $\beta$ be ordinals. We define the sum of $\alpha$ and $\beta$, denoted $\alpha+\beta$, to be the ordinal that is order isomorphic to the set $\alpha\sqcup\beta$ equipped with the order $\preceq$ given by $\eta\preceq\gamma$ if and only if either
@@
Observe that $\alpha+1$ is the successor of $\alpha$ for any ordinal $\alpha$. Note that addition of ordinals is not commutative, e.g., $11+\omega\neq\omega+$ 1, as $1+\omega$ is a limit ordinal (namely $\omega$) and $\omega+1$ is a successor ordinal. However, ordinal addition is associative.
@@
We define the product of $\alpha$ and $\beta$, denoted $\alpha\cdot\beta$, to be the ordinal that is order isomorphic to the set $\beta\times\alpha$ equipped with the lexicographical ordering.
@@
We define the exponentiation of $\alpha$ by $\beta$, denoted $\alpha^{\beta}$, to be the ordinal that is order isomorphic to the set of functions
$\{f\colon\thinspace\beta\to\alpha:|\{\eta\in\beta:f(\eta)\neq 0\}|<\infty\}$
@@ 2.3 Topology of ordinals
An ordinal has a canonical topology given by the order topology; for all that follows, whenever we reference the topology of an ordinal, we are referring to the order topology. Let $\alpha$ be an ordinal. A subbasis for the order topology on $\alpha$ is given by the sets of form $[0,\beta)=\{\eta\in\alpha:\eta<\beta\}$ and $(\beta,\alpha)=\{\eta\in\alpha:\beta<\eta\}$, where $\beta\in\alpha$.
Now, let $\lambda\in\alpha$ be a limit ordinal. Note that if $\beta<\lambda$ then $\beta+1<\lambda$, and observe that if $\lambda\in\alpha$ is a limit ordinal, then $[0,\lambda)=\bigcup_{\eta<\lambda}[0,\eta+1)$. It follows that the sets of the form $[0,\beta+1)=[0,\beta]$ and $(\beta,\alpha)=[\beta+1,\alpha)$ with $\beta\in\alpha$ form a subbasis for the order topology. In particular, the order topology admits a basis consisting of clopen sets, i.e., it is zero-dimensional. Moreover, it is an exercise to check that sets of the form $[\beta,\gamma]$ are compact and that if $\lambda$ is a limit ordinal then $[0,\lambda)$ fails to be compact. We summarize this in the following proposition.
@@
For the remainder of the article, we will focus on successor ordinals, and hence, the compact ordinals. The next goal of this subsection is to classify the successor ordinals up to homeomorphism, which is accomplished in 2.15. In the process, we establish several preliminary results and introduce the Cantor–Bendixson derivative.
@@
By definition, $\alpha+\beta$ is order isomorphic to the disjoint union of $\alpha$ and $\beta$ with the order $\preceq$ given in Definition 2.3; call this well-ordered set $X$ and equip it with its order topology, so that $X$ is homeomorphic to $\alpha+\beta$. The inclusions of $\alpha$ and $\beta$ into $X$ are topological embeddings. The image of $\beta$ in $X$ is closed and the image of $\alpha$ is open. As $\alpha$ is compact, it must also be closed. Therefore, both $\alpha$ and $\beta$ are clopen subsets of $X$, implying $X$, and hence, $\alpha+\beta$ is homeomorphic to the topological disjoint union of $\alpha$ and $\beta$. Similarly, $\beta+\alpha$ is homeomorphic to the topological disjoint union of $\alpha$ and $\beta$, implying that $\alpha+\beta$ and $\beta+\alpha$ are homeomorphic. ∎
@@
As the limit capacity of $\beta$ is strictly less than $\alpha$, it follows that $\beta\cdot\omega\leq\alpha$, as $\beta+k<\alpha$ for all $k\in\omega$. Let $\gamma$ be the ordinal order isomorphic to the complement of $\beta\cdot\omega$ in $\alpha$. Then $\alpha=\beta\cdot\omega+\gamma$. So, $\beta+\alpha=\beta+(\beta\cdot\omega+\gamma)=(\beta+\beta\cdot\omega)+\gamma$. It is readily checked that $\beta+\beta\cdot\omega=\beta\cdot\omega$, and hence $\beta+\alpha=\alpha$. ∎
@@
Let $\alpha$ be an ordinal. If $\beta$ is the limit capacity of $\alpha$ and $k$ is its coefficient, then $\alpha$ is homeomorphic to $\omega^{\beta}\cdot k+1$.
@@
where $\gamma$ is a successor ordinal. By 2.9, $\alpha$ is homeomorphic to $\gamma+\left(\omega^{\beta}\cdot k+1\right)$. As the limit capacity of $\gamma$ is strictly less than that of $\omega^{\beta}\cdot k+1$, 2.10 implies that $\gamma+\left(\omega^{\beta}\cdot k+1\right)=\omega^{\beta}\cdot k+1$, and hence $\alpha$ is homeomorphic to $\omega^{\beta}\cdot k+1$. ∎
@@
Next, we introduce the Cantor–Bendixson degree and rank. We will not make the most general definition, rather focusing on a definition that suits our purposes.
Definition 2.13 (Cantor–Bendixson rank and degree).
Let $X$ be a topological space. The Cantor–Bendixson rank of $X$ is the least ordinal $\alpha$ such that $X^{(\alpha)}$ is empty, if such an ordinal exists; otherwise, it is set to be $\infty$. If $X$ is compact and its Cantor–Bendixson degree is an ordinal $\alpha$, then $\alpha$ is necessarily a successor ordinal, allowing us to define the Cantor–Bendixson degree of $X$ to be the cardinality of $X^{(\alpha-1)}$, which is finite.
Using the Cantor normal form and transfinite induction, it is an exercise to show the following:
@@
The classification of successor ordinals up to homeomorphism now follows immediately from 2.11 and 2.14.
@@
Let $X$ be a topological space of Cantor–Bendixson rank $\alpha$, with $\alpha\neq\infty$. For $x\in X$, we define the rank of $x$ to be the least ordinal $\beta$ such that $x\notin X^{(\beta)}$.
@@
Observe that if $\eta\leq\beta$, then $U$ contains an element of rank $\eta$. It follows that, as $U$ is a well-ordered set, $U$ must be isomorphic to an ordinal of Cantor–Bendixson rank $\beta+1$ and Cantor–Bendixson degree 1. Therefore, by 2.15, $U$ homeomorphic to $\omega^{\beta}+1$. ∎
Note that all sufficiently small neighborhoods of $b$ in $\alpha$ satisfy the hypothesis of 2.18, allowing us to conclude that $b$ is stable.
@@
As already noted, 2.18 says that all sufficiently small clopen neighborhoods of $b$ in $\alpha$ are homeomorphic. It will be useful to have a name for such sets.
@@
For an ordinal $\alpha$, a clopen neighborhood $U$ of an element $b$ in a $\alpha+1$ is stable if $b$ is the unique highest rank element in $U$.
By 2.18, a stable neighborhood of a rank $\beta+1$ element of $\alpha+1$ is homeomorphic to $\omega^{\beta}+1$.
@@
For any ordinal $\alpha$, the spaces $\omega^{\alpha+1}$ and ×$\mathbb{N}\times(\omega^{\alpha}+1)$ are homeomorphic.
@@
For each $k\in\mathbb{N}$, $\omega^{\alpha}\cdot k+1\in\omega^{\alpha+1}$. Now, for $\eta\in\omega^{\alpha+1}$, there exists $k\in\mathbb{N}$ such that $\eta\in\omega^{\alpha}\cdot k+1$. It follows that
$\omega^{\alpha+1}$ $=\bigcup_{k\in\mathbb{N}}(\omega^{\alpha}\cdot k+1)$
$=[0,\omega^{\alpha}]\cup\bigcup_{k\in\mathbb{N}}(\omega^{\alpha}\cdot k,\omega^{\alpha}\cdot(k+1)]$
$=[0,\omega^{\alpha}]\cup\bigcup_{k\in\mathbb{N}}[\omega^{\alpha}\cdot k+1,\omega^{\alpha}\cdot(k+1)]$
@@
where we let $I_{0}=[0,\omega^{\alpha}]$ and $I_{k}=[\omega^{\alpha}\cdot k+1,\omega^{\alpha}\cdot(k+1)]$ for $k\in\mathbb{N}$. Now, for distinct $j,k\in\omega$, $I_{k}$ is clopen and $I_{k}\cap I_{j}=\varnothing$. Therefore, $\omega^{\alpha+1}$ is homeomorphic to the topological disjoint union of the $I_{k}$. To finish, observe that each of the $I_{k}$ is order isomorphic to, and hence homeomorphic to, $\omega^{\alpha}+1$, implying that $\omega^{\alpha+1}$ is homeomorphic to $\omega\times(\omega^{\alpha}+1)$, which is itself is clearly homeomorphic to ×$\mathbb{N}\times(\omega^{\alpha}+1)$. ∎

### v2

@@ 1 Introduction
Let us finish with exploring how our work has implications for mapping class groups. For a countable ordinal $\alpha$, let us identify $X_{\alpha,1}$ with a subset of $\mathbb{S}^{2}$. Our first slate of results fail for $\operatorname{MCG}(\mathbb{S}^{2}\smallsetminus X_{\alpha,1})$. Let us explain: Using the work of Domat–Dickmann [10], Malestein–Tao [21] showed that $\operatorname{MCG}(\mathbb{S}^{2}\smallsetminus X_{0,1})$ surjects onto the additive group of real numbers. Now, there is an epimorphism $\operatorname{MCG}(\mathbb{S}^{2}\smallsetminus X_{\alpha,1})\to\operatorname{MCG}(\mathbb{S}^{2}\smallsetminus X_{0,1})$ obtained by “filling in” the ends whose capacity is less than $\alpha$ (this is called a forgetful homomorphism in the literature). It follows that $\operatorname{MCG}(\mathbb{S}^{2}\smallsetminus X_{\alpha,1})$ also surjects onto the reals; in particular, $\operatorname{MCG}(\mathbb{S}^{2}\smallsetminus X_{\alpha,1})$ is neither uniformly perfect, finitely normally generated, strongly bounded, nor strongly distorted. Therefore, our results together with the work of Malestein–Tao, show that the failure of $\operatorname{Homeo}(\mathbb{S}^{2}\smallsetminus X_{\alpha,1})$ to be strongly bounded, uniformly perfect, and finitely normally generated, stems from a two-dimensional obstruction, i.e., it cannot be detected by the action on the end space. It is also interesting to note that the results only fail in the category of abstract groups, meaning their analogs in the category of topological groups hold; in particular, Mann–Rafi [22] showed that $\operatorname{MCG}(\mathbb{S}^{2}\smallsetminus X_{\alpha,1})$ is coarsely bounded, and it follows from the work of the last author with Lanier [18] that $\operatorname{MCG}(\mathbb{S}^{2}\smallsetminus X_{\alpha,1})$ is topologically normally generated by a single mapping class (Baik [3] also showed finite topological normal generation for $\operatorname{MCG}(\mathbb{S}^{2}\smallsetminus X_{\alpha,d})$, where $\alpha$ is a countable ordinal and $d\in\mathbb{N}$).
@@ Acknowledgments
The authors thank Kathryn Mann for the reference to Galvin’s work and Ferrán Valdez for a discussion of his forthcoming work with collaborators on the geometry of homeomorphism groups of ordinals. We also thank the referee who provided a number of invaluable comments that have improved our exposition immensely.
@@ 2 Ordinals and their topology
We provide an introduction to ordinals and their topology, aimed at non-experts. In the first three subsections we cover the basics of ordinals, see for instance [20, Chapter 2], [9, §4.1–2], and [16, Chapter 2] for the set-theoretic aspects.
@@ 2.1 Defining the ordinals
A binary relation $\leq$ on a set $X$ is a (strict) total order if:
  • it is reflexive: $x\leq x$,
  • it is transitive: if $x\leq y$ and $y\leq z$ then $x\leq z$,
  • it is antisymmetric: if $x\leq y$ and $y\leq x$ then $x=y$, and
  • it is strongly connected: $x\leq y$ or $y\leq x$
for all $x,y,z\in X$. If, in addition, for any non-empty subset $A$ of $X$ there exists $a\in A$ such that $a\leq b$ for all $b\in A$, then $\leq$ is well-ordered. A well-ordered set is a pair $(X,\leq)$. Two well-ordered sets $X$ and $Y$ are order isomorphic if there exists an order-preserving bijection $X\to Y$. The strict total order $\leq$ gives rise to a non-strict order <<, which is defined by $x<y$ if and only if $x\leq y$ and $x\neq y$.
@@
An ordinal is the order-isomorphism class of a well-ordered set.
This definition has the advantage of being easy to state and convenient for introducing ordinal arithmetic. However, in general, it will be immensely helpful to have a canonical representative of an ordinal, which we now describe.
We will use von Neumann’s definition of the ordinals.
Definition 2.2 (von Neumann ordinal).
A set $\alpha$ is an von Neumann ordinal if it satisfies the following:
@@
Every element of a von Neumann ordinal is itself a von Neumann ordinal. A von Neumann ordinal can be well-ordered by set inclusion, i.e., if $\alpha$ is a von Neumann ordinal and $\beta,\gamma\in\alpha$, declaring $\beta\leq\gamma$ if and only if $\beta\subseteq\gamma$ yields a well-ordering relation on $\alpha$. Note that if $\beta,\gamma\in\alpha$, then $\gamma<\beta$ if and only if $\gamma\in\beta$. In particular, if $\beta\in\alpha$, then $\beta=\{\eta\in\alpha:\eta<\beta\}$. If $\alpha$ and $\beta$ are order-isomorphic von Neumann ordinals, then $\alpha=\beta$; it now follows from basic properties of well-ordered sets that given two von Neumann ordinals $\alpha$ and $\beta$, either $\alpha=\beta$, $\alpha\in\beta$, or $\beta\in\alpha$. We can therefore compare any two distinct von Neumann ordinals, writing $\beta<\alpha$ if and only if $\beta\in\alpha$ (here, we have slightly abused notation, using << to denote both a relation on a given von Neumann ordinal and as a comparison of two distinct von Neumann ordinals; however, no confusion arises with this abuse). In other words,
The von Neumann ordinal $\alpha$ is the set of all von Neumann ordinals less than $\alpha$.
Now, every well-ordered set is order-isomorphic to a unique von Neumann ordinal, implying that a von Neumann ordinal is a canonical representative of its order-isomorphism class.
Notation. Throughout the paper, we will abuse notation and conflate an ordinal with its associated von Neumann ordinal, using the word ordinal to refer to both simultaneously. It should clear from context whether we are using an ordinal as an isomorphism class or working directly with its von Neumann representative. We will write ordinals using lower-case Greek letters.
Given any ordinal $\alpha$, the set $\alpha\cup\{\alpha\}$ is an ordinal, called the successor of $\alpha$. An ordinal is a successor ordinal if it is the successor to some ordinal; otherwise, it is a limit ordinal [^2] We follow [9] in considering 0 a limit ordinal.. From the discussion above, any set of ordinals has a least element; this least element is necessarily the empty set and is denoted by 0. Therefore, 0 is the first limit ordinal. We can now proceed to construct the natural numbers, identifying 1 with $\{0\}$, or equivalently, $\{\varnothing\}$, 2 with $\{0,1\}$, or equivalently, $\{\varnothing,\{\varnothing\}\}$, etc. Throughout, we will identity the natural numbers, denoted $\mathbb{N}$, with $\mathbb{N}=\{1,2,\ldots\}$; in particular, $0\notin\mathbb{N}$. The natural numbers together with 0 constitute the set of finite ordinals, which itself is an ordinal, denoted $\omega$; equivalently, $\omega$ is the first infinite ordinal. Note that $\omega$ is the first non-zero limit ordinal.
@@
Let $\alpha$ be an ordinal, and let $P(\beta)$ be a property defined for all $\beta\leq\alpha$. If
  1. (i)
    (base case) $P(0)$,
  2. (ii)
    (successor case) $P(\beta)$ implies $P(\beta+1)$, and
  3. (iii)
    (limit case) for a limit ordinal $\lambda$, $P(\beta)$ for all $\beta<\lambda$ implies $P(\lambda)$.
then $P(\alpha)$. ∎
@@ 2.2 Ordinal arithmetic
We now turn to defining ordinal arithmetic, including the notions of addition, multiplication, and exponentiation.
@@
Let $\alpha$ and $\beta$ be ordinals. We define the sum of $\alpha$ and $\beta$, denoted $\alpha+\beta$, to be the order-isomorphism class of the set $\alpha\sqcup\beta$ equipped with the total order that extends the orders on $\alpha$ and $\beta$ and declares the elements of $\alpha$ to be smaller than those of $\beta$, that is, $\alpha\sqcup\beta$ equipped with the order $\preceq$ given by $\eta\preceq\gamma$ if and only if either
@@
Observe that $\alpha+1$ is the successor of $\alpha$ for any ordinal $\alpha$. Note that addition of ordinals is not commutative, e.g., $1+\omega\neq\omega+1$, as $1+\omega$ is a limit ordinal (namely $\omega$) and $\omega+1$ is a successor ordinal. However, ordinal addition is associative.
@@
We define the product of $\alpha$ and $\beta$, denoted $\alpha\cdot\beta$, to be the order-isomorphism class of the set $\beta\times\alpha$ equipped with the lexicographical ordering.
@@
Though commutativity of addition and multiplication both fail, it is readily verified that associativity and left-distributivity hold. We record this fact in the following lemma.
Lemma 2.6.
Let $\alpha,\beta$, and $\gamma$ be ordinals. Then
$(\alpha+\beta)+\gamma=\alpha+(\beta+\gamma)$
and
$\alpha\cdot(\beta+\gamma)=\alpha\cdot\beta+\alpha\cdot\gamma.$
@@
We define the exponentiation of $\alpha$ by $\beta$, denoted $\alpha^{\beta}$, to be the ordinal that is order isomorphic to the set of finitely supported functions
$\{f\colon\thinspace\beta\to\alpha:f(\eta)=0\text{ for all but finitely many }\eta\}$
@@ 2.3 Topology of ordinals
An ordinal has a canonical topology given by the order topology; for all that follows, whenever we reference the topology of an ordinal, we are referring to the order topology. Let $\alpha$ be an ordinal. A subbasis for the order topology on $\alpha$ is given by the sets of form $0,\beta)=\{\eta\in\alpha:\eta<\beta\}$ and $(\beta,\alpha)=\{\eta\in\alpha:\beta<\eta\}$, where $\beta\in\alpha$.
Now, let $\lambda\in\alpha$ be a limit ordinal. Note that if $\beta<\lambda$ then $\beta+1<\lambda$, and observe that if $\lambda\in\alpha$ is a limit ordinal, then $0,\lambda)=\bigcup_{\eta<\lambda}[0,\eta+1)$. It follows that the sets of the form $0,\beta+1)=[0,\beta]$ and $(\beta,\alpha)=[\beta+1,\alpha)$ with $\beta\in\alpha$ form a subbasis for the order topology. In particular, the order topology admits a basis consisting of clopen sets, i.e., it is zero-dimensional. Moreover, it is an exercise to check that sets of the form $\beta,\gamma]$ are compact and that if $\lambda$ is a limit ordinal then $0,\lambda)$ fails to be compact. We summarize this in the following proposition.
@@
The next goal of this subsection is to classify the successor ordinals up to homeomorphism, which is accomplished in 2.18. In the process, we establish several preliminary results and introduce the Cantor–Bendixson derivative.
@@
By definition, the ordinal $\alpha+\beta$ is represented by the disjoint union of $\alpha$ and $\beta$ with the order $\preceq$ given in Definition 2.4. In the space $\alpha+\beta$, $\beta$ is closed and $\alpha$ is open. As $\alpha$ is compact, it must also be closed. Therefore, both $\alpha$ and $\beta$ are clopen subsets of $\alpha+\beta$, implying $\alpha+\beta$ is topologically the disjoint union of $\alpha$ and $\beta$. Similarly, $\beta+\alpha$ is homeomorphic to the topological disjoint union of $\alpha$ and $\beta$, implying that $\alpha+\beta$ and $\beta+\alpha$ are homeomorphic. ∎
@@
As the limit capacity of $\beta$ is strictly less than $\alpha$, it follows that $\beta\cdot\omega\leq\alpha$, as $\beta+k<\alpha$ for all $k\in\omega$. Let $\gamma$ be the ordinal of the interval $(\beta\cdot\omega,\alpha)$ in $\alpha$. Then $\alpha=\beta\cdot\omega+\gamma$. Using associativity and distributivity (2.6), we have:
$\beta+\alpha=\beta+(\beta\cdot\omega+\gamma)=(\beta+\beta\cdot\omega)+\gamma=\beta\cdot(1+\omega)+\gamma=\beta\cdot\omega+\gamma=\alpha$
as desired. ∎
@@
Let $\alpha$ be a successor ordinal. If $\beta$ is the limit capacity of $\alpha$ and $k$ is its coefficient, then $\alpha$ is homeomorphic to $\omega^{\beta}\cdot k+1$.
@@
where $\gamma$ is a successor ordinal. By 2.11, $\alpha$ is homeomorphic to $\gamma+\left(\omega^{\beta}\cdot k+1\right)$. As the limit capacity of $\gamma$ is strictly less than that of $\omega^{\beta}\cdot k+1$, 2.12 implies that $\gamma+\left(\omega^{\beta}\cdot k+1\right)=\omega^{\beta}\cdot k+1$, and hence $\alpha$ is homeomorphic to $\omega^{\beta}\cdot k+1$. ∎
@@
Next, we introduce the Cantor–Bendixson rank and degree.
Definition 2.15 (Cantor–Bendixson rank).
Let $X$ be a topological space. The Cantor–Bendixson rank of $X$ is the least ordinal $\alpha$ such that $X^{(\alpha)}$ is perfect.
As the intersection of nonempty nested closed sets is nonempty in a compact space, if $X$ is a compact space whose Cantor–Bendixson derivatives stabilize in the empty set, then its Cantor–Bendixson rank is necessarily a successor ordinal and the last non-empty Cantor–Bendixson derivative has finite cardinality, allowing us to define the Cantor–Bendixson degree.
Definition 2.16 (Cantor–Bendixson degree).
Let $X$ be a compact topological space whose Cantor–Bendixson derivatives are eventually empty. If $\alpha+1$ is the Cantor–Bendixson rank of $X$, then the Cantor–Bendixson degree of $X$ is the cardinality of $X^{(\alpha)}$.
A successor ordinal is compact, non-perfect, and its Cantor–Bendixson derivatives are eventually empty. Therefore, we can investigate both is Cantor–Bendixson rank and degree. Using the Cantor normal form and transfinite induction, it is an exercise to show the following:
@@
The classification of successor ordinals up to homeomorphism now follows immediately from 2.13 and 2.17.
@@
Let $X$ be a topological space of Cantor–Bendixson rank $\alpha$ satisfying $X^{(\alpha)}=\varnothing$. For $x\in X$, we define the rank of $x$ to be the least ordinal $\beta$ such that $x\notin X^{(\beta)}$.
Observe that the rank of a point is necessarily a successor ordinal. By 2.17, the rank of an element in an ordinal is the successor to its capacity. Despite discussing the capacity of a point in the introduction, we will generally discuss the rank going forward, as it is topologically defined.
@@
Observe that if $\eta\leq\beta$, then $U$ contains an element of rank $\eta$. It follows that, as $U$ is a well-ordered set, $U$ must be isomorphic to an ordinal of Cantor–Bendixson rank $\beta+1$ and Cantor–Bendixson degree 1. Therefore, by 2.18, $U$ homeomorphic to $\omega^{\beta}+1$. ∎
Note that all sufficiently small neighborhoods of $b$ in $\alpha$ satisfy the hypothesis of 2.21, allowing us to conclude that $b$ is stable.
@@
As already noted, 2.21 says that all sufficiently small clopen neighborhoods of $b$ in $\alpha$ are homeomorphic. It will be useful to have a name for such sets.
@@
For an ordinal $\alpha$ and $b\in\alpha+1$, a clopen neighborhood $U$ of $b$ in $\alpha+1$ is stable if $b$ is the unique highest rank element in $U$.
By 2.21, a stable neighborhood of a rank $\beta+1$ element of $\alpha+1$ is homeomorphic to $\omega^{\beta}+1$.
@@
Let $\alpha$ be an ordinal. If $X$ is a countable discrete space, then the spaces $\omega^{\alpha+1}$ and $X\times(\omega^{\alpha}+1)$ are homeomorphic.
@@
For each $k\in\omega$, $\omega^{\alpha}\cdot k+1\in\omega^{\alpha+1}$. Now, for $\eta\in\omega^{\alpha+1}$, there exists $k\in\mathbb{N}$ such that $\eta\in\omega^{\alpha}\cdot k+1$. It follows that
$\omega^{\alpha+1}$ $=\bigcup_{k\in\mathbb{N}}[0,\omega^{\alpha}\cdot k]$
$=\{0\}\cup\bigcup_{k\in\omega}[\omega^{\alpha}\cdot k+1,\omega^{\alpha}\cdot(k+1)]$
@@
where we let $I_{0}=[0,\omega^{\alpha}]$ and $I_{k}=[\omega^{\alpha}\cdot k+1,\omega^{\alpha}\cdot(k+1)]$ for $k\in\mathbb{N}$. Now, for distinct $j,k\in\omega$, $I_{k}$ is clopen and $I_{k}\cap I_{j}=\varnothing$. Therefore, $\omega^{\alpha+1}$ is homeomorphic to the topological disjoint union of the $I_{k}$. To finish, observe that each of the $I_{k}$ is order isomorphic to, and hence homeomorphic to, $\omega^{\alpha}+1$, implying that $\omega^{\alpha+1}$ is homeomorphic to $\omega\times(\omega^{\alpha}+1)$. As as $\omega$ is countable and discrete, it is homeomorphic to $X$, implying $\omega\times(\omega^{\alpha}+1)$, and hence $\omega^{\alpha+1}$, is homeomorphic to $X\times(\omega^{\alpha}+1)$. ∎